OpenStudy (anonymous):

A circle has a diameter of ten yards. What is its area? Use three point one four for pi

OpenStudy (anonymous):

@ganymede

OpenStudy (strawberry17):

Well, the formula for area is: A = πr^2 π = 3.14 The radius equal half of the diameter. So, r = 10/2 = 5 Now plug into Area formula: A = 3.14(5)^2 Make sure to first square 5 and then multiply that number by 3.14 and you will have your area :)